Let 0<θ<2π, θ ≠ π. Consider the linear transformation T: C^2→C^2 given by matrix

[ cosθ -sinθ](w.r.t standard basis)

[ sinθ cosθ]. Find the vector v1, v2 such that Tv1= e^iθv1, Tv2= e^-iθv2. Is {v1,v2} a basis for C^2? Give reason for your answer.


1
Expert's answer
2021-02-28T17:28:03-0500

Matrix of transformation:

"A=\begin{bmatrix}
 cos\theta & -sin\theta \\
 sin\theta & cos\theta
\end{bmatrix}"

Let

"v_1=\begin{bmatrix}
 x_1 \\
 x_2 
\end{bmatrix}" , "v_2=\begin{bmatrix}
 y_1 \\
 y_2 
\end{bmatrix}"

Then:

"T(v_1)=Av_1=\begin{bmatrix}
 cos\theta & -sin\theta \\
 sin\theta & cos\theta
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 
\end{bmatrix}"

"T(v_1)=\begin{bmatrix}
 x_1cos\theta-x_2sin\theta \\
 x_1sin\theta+x_2cos\theta 
\end{bmatrix}=\begin{bmatrix}
 x_1e^{i\theta} \\
 x_2e^{i\theta} 
\end{bmatrix}"


"x_1cos\theta-x_2sin\theta=x_1e^{i\theta}=x_1cos\theta+ix_1sin\theta"

"x_1sin\theta-x_2cos\theta=x_2e^{i\theta}=x_2cos\theta+ix_2sin\theta"


"-x_2=ix_1\implies x_1=-x_2/i=ix_2"

"x_2cos\theta=0"


"v_1=\begin{bmatrix}
 x_1 \\
 -ix_1
\end{bmatrix}=x_1\begin{bmatrix}
 1 \\
 -i
\end{bmatrix}"


"T(v_2)=Av_2=\begin{bmatrix}
 cos\theta & -sin\theta \\
 sin\theta & cos\theta
\end{bmatrix}\begin{bmatrix}
 y_1 \\
 y_2 
\end{bmatrix}"

"T(v_2)=\begin{bmatrix}
 y_1cos\theta-y_2sin\theta \\
 y_1sin\theta+y_2cos\theta 
\end{bmatrix}=\begin{bmatrix}
 y_1e^{-i\theta} \\
 y_2e^{-i\theta} 
\end{bmatrix}"


"y_1cos\theta-y_2sin\theta=y_1e^{-i\theta}=y_1cos\theta-iy_1sin\theta"

"y_1sin\theta-y_2cos\theta=y_2e^{-i\theta}=y_2cos\theta-iy_2sin\theta"


"y_2=iy_1"


"v_2=\begin{bmatrix}
 y_1 \\
 iy_1
\end{bmatrix}=y_1\begin{bmatrix}
 1 \\
 i
\end{bmatrix}"


The set "\{v_1,v_2\}" is lineary independent, and dimension of "\{v_1,v_2\}" is the same as dimension of "C^2" . So "\{v_1,v_2\}" is a basis for "C^2."